The sum of the ages of a woman and her son is 36 years. Six years from now the woman will be twice as old as her son. How old is each now?
Answer by TimothyLamb(4379)

---
x = woman
y = son
---
x + y = 36
x + 6 = 2(y + 6)
---
put the system of linear equations into standard form:
---
x + y = 36
x + 6 = 2y + 12
---
x + y = 36
x - 2y = 6
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = woman = 26
y = son = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php